OpenStudy (anonymous):

Solve, or determine the roots of the following quadratic equations by factoring. a) 3x^2+15x=0 b)x^2-64=0 c)x^2-6x-27=0

OpenStudy (anonymous):

a. x=0,-5

OpenStudy (anonymous):

b. x=8,-8

OpenStudy (anonymous):

c. x=9,-3

OpenStudy (anonymous):

How'd you get that ?

OpenStudy (anonymous):

a) x(3x+15)=0, so x=0 ; 3x+15=0; x=-5 b) x^2=64; x=sqrt(64); x=+8, -8 c) quadratic equation so; x=(-b+-sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

a. 3x(x-5)=0 => x=0,5 others are similar

OpenStudy (anonymous):

Oh alright, thanks !